Find the volume of a cube that has a total area of 96 sq. in.
Answer by ewatrrr(24785)


Hi
cube with 6 sides having total area of 96 in^2
96in^2/6 = 16in^2 per side s*s = Area (side) where s(length of a side) = 4in
Volume(cube) = s*s*s = 4in * 4in * 4in = 64 in^3